The side of a square house is 24 feet long and the house is located on a lot which is 50 feet longer than it is wide. The perimeter of the lot is 20 feet more than five times the perimeter of the house. find the length of the lot.
Answer by fractalier(6550)

If the house is square and has a side of 24 feet, its perimeter must be 96 feet.
For the lot, we have perimeter
P = 2L + 2W = 5(96) + 20
and
L = 50 + W
Now substitute this into the first equation and get
2(50+W) + 2W = 5(96) + 20
100 + 2W + 2W = 480 + 20 = 500
Subtract 100 and get
4W = 400
W = 100 feet
L = 50+W = 150 feet